As I really think your comment is very helpful, I was wondering whether you could help me with my current challenge:
Demand: Q = 200 – 5P.
Government protection: 20% tax on imported automobiles (your only competitors) and a $10 subsidy for every unit sold. It is assumed that every unit produced will sell, though the price will adjust until the point where the market will clear.
The cost curve for your organization is TC = 20 + 2Q.

How can I derive a supply function from these inputs? Is there a way?
To maximise revenue is it enough to just say MR = 40 - 0.4Q = 0?
What is the net effect of the subsidy provided by the government at the profit-maximising output?
1
Expert's answer
2016-08-09T07:17:02-0400
Answer on Question #61195, Economics / Microeconomics
Demand: Q = 200 – 5P, TC = 20 + 2Q, 20% tax on imported automobiles (your only competitors) and a $10 subsidy for every unit sold.
A supply function is the part of marginal cost curve MC after the intersection with AVC curve, so:
S = MC = TC' = 2.
The revenue is maximisedm when MR = 40 - 0.4Q = 0.
Profit is maximized, when MR = MC = P = $2, so 40 - 0.4Q = 2, Q = 95 units.
The net effect of the subsidy provided by the government at the profit-maximising output will cause the existance of the deadweight loss and decrease of the total welfare.